Consider the following data in Billions of dollars:

GDP 7500

Investment 1300

Transfer Payment 1050

Government Expenditures 1200

Exports 1050

Imports 950

Net Factor Payment 20






a) Use the above data to calculate the consumption.

b) Whether the economy has trade deficit or trade surplus? Why?


1
Expert's answer
2022-02-01T10:14:28-0500

Qn. A

Aggregate income, for the economy, assumed t equal to the GDP is given by;

"GDP=Y"

"Y=C+I+G+(X-M)+TP+Nfp"

"C= consumption"

"I=investment"

"G=Government\ expenditure"

"X-M=Net\ Exports"

"TP= Transfer\ Payment"

"Nfp=Net\ Factor\ payment"


"7500=C+1300+1200+(950-1050)+1050+20"

"7500=C+3470"

"\bold{Consumption=\$4030} Billion"


Qn. B

The economy has a trade surplus. A surplus in trade is seen where a positive balance is obtained from subtracting the total amount of imports from total exports. "1050-950=100"